OpenStudy (anonymous):

state the gradient and the coordinates of the y intercept for the following straight lines. y=3x-5

Directrix (directrix):

I assume that gradient means slope. y = mx + b is the slope and y-intercept form of a line. m is the slope (gradient) and b is the y-intercept.

Directrix (directrix):

y = mx + b y=3x-5 Compare. m = 3 in your problem b, the y-intercept = -5. You can write it as the ordered pair (0 , -5) if you wish.

OpenStudy (anonymous):

thanq

Directrix (directrix):

Glad to help.